OpenStudy (anonymous):

The base of a triangle is 4cm greater than the height. The area is 48cm^2. Find height and the length of the base. the height of the triangle is the base of the triangle is

OpenStudy (anonymous):

let x be the height then the base is x+4 S=x(x+4)/2=48 x^2+4x-96=0 solve to get:x=8 then height =8 cm base =12 cm

OpenStudy (mathteacher1729):

Area of a triangle = 1/2 * b * h h = h b = h + 4 So instead of A = 1/2 * b * h you have A = 1/2 * (h+4) * h And we're given that A = 48 so 48 = 1/2 * (h+4) * h Which is now a quadratic in h. Can you go figure it out from here?

OpenStudy (anonymous):

ty for the formula